OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x – 3y = –12

OpenStudy (anonymous):

set x=0 to get y-int 2(0)-3y=-12 y=4 (0,-4)

OpenStudy (anonymous):

Set y=0 to get x-int 2x-3(0)=-12 2x=-12 x=-6 (-6,0)